write a quadratic equation whose zeroes are 5,-5
Answer by Boreal(15235)

the zeros mean x=5 and x= -5
rewrite as (x-5)=0 and (x+5)=0
the factors are (x-5)(x+5)
an equation is x^2-25=0.